# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

### v1

@@ Abstract.
Applications of Large Language Models (LLMs) are rapidly growing in industry and academia for various software engineering (SE) tasks. As these models become more integral to critical processes, ensuring their reliability and trustworthiness becomes essential. Consequently, the concept of trust in these systems is becoming increasingly critical. Well-calibrated trust is important, as excessive trust can lead to security vulnerabilities, and risks, while insufficient trust can hinder innovation. However, the landscape of trust-related concepts in LLMs in SE is relatively unclear, with concepts such as trust, distrust, and trustworthiness lacking clear conceptualizations in the SE community. To bring clarity to the current research status and identify opportunities for future work, we conducted a comprehensive review of $88$ papers: a systematic literature review of $18$ papers focused on LLMs in SE, complemented by an analysis of $70$ papers from broader trust literature. Additionally, we conducted a survey study with $25$ domain experts to gain insights into practitioners’ understanding of trust and identify gaps between existing literature and developers’ perceptions. The result of our analysis serves as a roadmap that covers trust-related concepts in LLMs in SE and highlights areas for future exploration.
@@ 1. Introduction
The use of LLMs in SE is becoming increasingly prevalent, with applications spanning bug fixing  (Tufano et al., 2018), defect prediction  (Wang et al., 2016), input generation  (Liu et al., 2017), and various other tasks  (Gu et al., 2018a; Guo et al., 2017; Corley et al., 2015; Liu et al., 2018; Watson et al., 2021; Gu et al., 2018b). These models are now widely adopted by software developers, researchers, and students, demonstrating their potential to enhance productivity and automate complex SE tasks.
@@
However, the successful integration of LLMs into SE workflows depends not only on their technical capabilities but also on how practitioners perceive and trust them (Roychoudhury et al., 2025). For example, misaligned trust between LLM and practitioner, either excessive (i.e., over-trust) or insufficient (i.e., under-trust), can significantly impact the effectiveness and security of these LLMs. Achieving the right alignment is crucial, as both extremes entail undesirable consequences: over-trust in LLMs may lead to security vulnerabilities, data integrity risks, and erroneous decision-making  (Thorne, 2024), while under-trust can hinder adoption, reducing the potential benefits of LLMs  (Balayn et al., 2024).
@@
The alignment of trust between LLMs and practitioners determines the myriad of feasible trust relations within software engineering settings. Trust relations are characterized not only by the trustor, defined as the party that decides whether or not to trust, and the trustee, defined as the party or the object to be trusted  (Harjit Sekhon and Devlin, 2014), but also by a specific activity  (Balayn et al., 2024). For example, a typical trust relation scenario is when a practitioner (i.e., trustor) accepts or rejects a snippet generated by an LLM such as GitHub Copilot (i.e., trustee). Furthermore, the perception of trust varies among the trustor, SE-tasks, and domains (e.g., SE, HCI, DL) within these relationships (Balayn et al., 2024). For example, a developer may confidently rely on an LLM to generate boilerplate code but hesitate to use it for test case generation due to concerns about domain-specific accuracy. Understanding the core concepts of trust in these contexts is crucial for uncovering the complexities of these relationships.
@@
However, core concepts such as trust, distrust, and trustworthiness are often used interchangeably, obscuring their precise meanings  (Sharma et al., 2024). As a result, fundamental questions about trust in LLM4SE remain unresolved: What factors shape trust, and do these factors vary across experience levels, such as between expert and novice developers, or across different SE tasks? Is distrust simply the absence of trust, or does it represent a distinct state? Addressing these unresolved questions is crucial for adopting LLMs in SE workflows successfully, as demonstrated by research findings in the field of Human-Computer Interaction (HCI) (Choung et al., 2022; Glikson and Woolley, 2020). These HCI studies have shown that trust is a crucial factor that influences the willingness of stakeholders to adopt LLMs. Unfortunately, existing studies provide a limited foundational understanding of how trust is defined, the factors that influence it, and the metrics used to assess it.
@@
To establish this foundational understanding, we must first assess the current state of research on trust in LLM4SE. Nonetheless, a comprehensive review of existing work has not been conducted, making it difficult to determine what aspects of trust have been explored and where significant gaps remain. The absence of a structured overview of trust concepts hinders efforts to define key trust-related concepts, identify influencing factors, and develop reliable evaluation methodologies. Additionally, the SE practitioners’ perspective on the definition, factors, and metrics of trust remains largely under-explored, leaving a disconnect between academic research and real-world needs. Without a clear understanding of the existing landscape, advancing trust-aware LLM systems for SE remains challenging.
@@
To address these challenges, we conducted a comprehensive approach that combines literature review with practitioner insights. First, we reviewed $18$ SE-focused papers to examine how trust-related concepts are defined, what factors influence trust in LLM4SE, and what metrics are used for evaluation. Second, we conducted a survey study with $25$ SE practitioners to capture real-world perspectives on trust in LLM4SE, identify task-specific trust factors, and explore variations in trust perceptions at different levels of expertise. In addition, we carried out a complementary analysis of $70$ papers from disciplines such as DL, HCI, and automation to enrich and complement the trust conceptualization from other fields.
@@
By integrating findings from the software engineering literature and practitioner experiences, we offer a comprehensive perspective on the current state of trust research in LLM4SE. Beyond identifying gaps in definitions, influencing factors, and evaluation methodologies, our study states the limitations of existing approaches and highlights the pressing need for standardized trust assessment frameworks. Additionally, we examine the disconnect between academic research and practitioner expectations, revealing key areas where trust calibration –the alignment of users’ trust level with actual reliability or trustworthiness of a system (Turner et al., 2024) – remains a challenge in real-world SE workflows. To improve these aspects, we propose concrete steps to validate trust definitions, improve trust factors, and establish trust metrics. By systematically synthesizing these findings, this work lays the groundwork for future advancements in trust-aware LLM integration, offering concrete directions for improving the reliability, interpretability, and user confidence in these systems.
@@
We structure our investigation around three key Research Questions (RQs) to map current understanding of trust concepts in LLM4SE. Through a systematic literature review and a survey study with SE practitioners, we examine how trust, distrust, and trustworthiness are defined in the context of LLM4SE (RQ1), identify the factors that shape trust perceptions (RQ2), and analyze existing approaches to measuring trust in LLMs (RQ3). Our findings reveal that while LLMs are increasingly adopted in SE, trust remains a loosely defined and inconsistently measured construct. Existing research neglects domain-specific definitions of trust-related concepts, often borrowing from other disciplines without adaptation to SE. In addition, trust factors, such as precision, interpretability, robustness, and workflow integration, are not consistently considered in different studies, leading to a fragmented understanding of the factors driving trust in LLM4SE. Finally, trust evaluation remains an open challenge, with limited efforts to establish standardized, multi-dimensional metrics that account for the complexity of trust in SE workflows.
@@
    We demonstrate a current understanding of trust definitions, factors that influence trust in LLM4SE, and the metrics used to evaluate trust in LLM4SE, through an extensive review of the literature, including $18$ SE-focused papers, and analysis $70$ papers from broader trust literature.
@@
    We complement our review with a survey study of $25$ participants to map current research and gather practitioners’ perspectives on trust in LLMs in SE. We highlight gaps between literature and practical scenarios and provide a roadmap for future research.
@@
    We propose precise definitions for trust, distrust, and trustworthiness in the context of LLM4SE rooted in the existing literature and perspectives of practitioners, identify key factors influencing trust in LLM4SE and group them into model-specific and user-centric factors. Furthermore, we highlight the limitations of single-item metrics and advocate for the development of comprehensive evaluation frameworks that better capture the complexity of trust in LLM4SE.
@@ 2. Why are Trust, Distrust, and Trustworthiness relevant for researching LLMs in Software Engineering?
Misplaced trust, whether excessive or insufficient, can lead to poor decisions and unintended consequences such as security vulnerabilities, code smells, and data leaks (Thorne, 2024; Mohsin et al., 2024; Balayn et al., 2024). For instance, researchers recently discovered that AI-powered code assistants, including GitHub Copilot and ChatGPT, often generate insecure code snippets that developers often failed to recognize as flawed (Sandoval et al., 2023; Pearce et al., 2025). The study also found that when LLM-generated suggestions contained subtle security vulnerabilities, developers were more likely to trust and use them without sufficient scrutiny, leading to increased security risks (Pearce et al., 2025). This tendency to over-rely on LLM-generated code without thorough verification goes beyond individual mistakes: it reflects a growing shift in how developers interact with LLMs and make critical coding decisions. As LLMs become integral to software development, they do more than suggest code; they influence workflows, assumptions, and even risk perception. The growing reliance on LLMs in SE requires a structured approach to understanding trust, distrust, and trustworthiness (Huang et al., 2024; Lo, 2023).
@@
The structured understanding of trust in LLM4SE can draw inspiration from research in other fields such as HCI and AI ethics. Studies in these fields highlight key factors that influence trust, including accuracy, interpretability, and transparency (Glikson and Woolley, 2020; Sharma et al., 2024). However, trust in SE is particularly complex due to the technical nature of the domain, the potential for hidden errors in generated code, and the need for alignment with best practices (Cui et al., 2024; Balayn et al., 2024). Unlike traditional software tools, LLMs produce non-deterministic outputs, meaning the same prompt may yield different responses, making trust assessment even more challenging (Zhu et al., 2024; Ye et al., 2023). Recognizing these challenges, researchers have called for a deeper investigation into trust in SE, e.g., D. Lo (Lo, 2023) emphasizes that as software development moves toward Software 2.0, where LLMs play a central role in coding, ensuring trustworthiness in AI-assisted development is imperative.
@@
While these calls to research highlight the urgency of understanding trust in LLM4SE, existing studies have yet to comprehensively address this issue. Despite the significance of trust in AI-powered SE tools, existing research on LLM4SE has largely overlooked definitions, factors, and metrics of trust. Our review identifies three key challenges related to trust in software engineering research. First ($CH_{1}$), only a small fraction of SE studies explicitly define trust, and distinctions between trust, distrust, and trustworthiness remain underexplored, making it difficult to assess their role in real-world SE workflows. Second ($CH_{2}$), while the broader trust literature provides valuable insight into the factors that influence trust, these perspectives have not been adequately integrated into SE research, leaving a gap in understanding what affects trust in developer interactions with LLMs. Third ($CH_{3}$), there is a noticeable absence of standardized metrics for measuring trust, and the perspectives of software practitioners on how trust should be quantified and evaluated are not well represented.
@@
To address these challenges, we conduct a structured literature review, analyze trust research beyond SE, and complement our findings with a practitioner survey. In doing so, our goal is to provide a comprehensive understanding of trust in LLM4SE and lay the groundwork for future research and the development of trustworthy LLM-powered coding assistants aligned with developers.
@@ 3. Research Questions
This study investigates trust in LLM4SE through three RQs. Each question is addressed using a systematic review of trust literature in LLM4SE and a survey with practitioners. To provide additional context, each research question is complemented with an analysis of the underlying trust concept from the broader trust literature.
@@ 1. RQ1
    Definition of Trust: How is trust in LLMs defined and conceptualized in the context of SE? This question aims to explore how trust, distrust, and trustworthiness are defined by analyzing existing LLM4SE literature and practitioner perspectives and address $CH_{1}$. These findings are contextualized through definitions found in broader trust literature to highlight key similarities, differences, and gaps. By analyzing all these perspectives, we provide refined, evidence-based definitions that align with academic and practitioner expectations.
@@ 2. RQ2
    Antecedents of Trust: What are the antecedents of trust (i.e. factors that may lead to trust) in LLM in the context of SE? This question investigates the key factors that lead to the formation of trust in LLM4SE, drawing from both LLM4SE literature and practitioner experiences and address $CH_{2}$. Additional context is provided by analyzing trust factors in established trust models and related domains. The analysis provides actionable recommendations for fostering trust in LLM4SE.
@@ 3. RQ3
    Trust Metrics: How is trust in LLMs measured in the context of SE? This question focuses on evaluating existing metrics used to assess trust in LLM4SE to address $CH_{3}$. We investigate existing trust assessment methods in LLM4SE, examine how professionals evaluate trust in real-world settings, and compare these with broader trust evaluation frameworks. By integrating these perspectives, we propose a road map for developing a comprehensive and domain-specific trust metric for LLM4SE.

### v2

@@ Abstract.
The application of Large Language Models (LLMs) in Software Engineering (SE) continues to grow rapidly across both industry and academia. As these models become integral to critical SE processes, ensuring their reliability and trustworthiness becomes essential. Achieving this requires a balanced approach to trust: excessive trust can introduce security vulnerabilities, while insufficient trust may hinder innovation. However, the conceptual landscape of trust in LLMs for SE(LLM4SE) remains unclear. Key concepts such as trust, distrust, and trustworthiness lack precise definitions, factors that shape trust formation remain underexplored, and metrics for trust in LLMs remain undeveloped. To clarify the current research landscape and identify future directions, we conducted a comprehensive review of $88$ articles: a systematic review of $18$ studies on LLMs in SE, supplemented by an analysis of $70$ articles from the broader trust literature. Furthermore, we surveyed $25$ domain experts to gather practitioners’ perspectives on trust and identify gaps between their experiences and the existing literature. Our findings provide a structured overview of trust-related concepts in LLM4SE, outlining key areas for future research. This study contributes to building more trustworthy LLM-assisted software engineering processes, ultimately supporting safer and more effective adoption of LLMs in SE.
@@ 1. Introduction
The use of LLMs in SE is becoming increasingly prevalent, with applications spanning bug fixing (Tufano et al., 2018), defect prediction (Wang et al., 2016), input generation (Liu et al., 2017), and various other tasks (Gu et al., 2018; Guo et al., 2017; Corley et al., 2015; Liu et al., 2018; Watson et al., 2021). These models are now widely adopted by software developers, researchers, and students, demonstrating their potential to enhance productivity and automate complex SE tasks.
@@
However, the successful integration of LLMs into SE workflows depends not only on their technical capabilities, but also on how practitioners perceive and trust them (Roychoudhury et al., 2025). For example, misaligned trust between the LLM and the practitioner, either excessive (i.e., overtrust) or insufficient (i.e., undertrust), can significantly impact the effectiveness and security of these LLMs. A recent empirical study on ChatGPT and Gemini in Java projects found that up to 7% of automatic refactorings broke functionality or introduced syntax errors(Liu et al., 2024a), demonstrating risks of overtrust. Conversely, Peng et al.’s controlled study found that developers using GitHub Copilot completed a JavaScript HTTP server task 55.8% faster(Peng et al., 2023), highlighting significant productivity gains achievable with appropriate trust. Hence, achieving the right alignment is crucial, as both extremes entail undesirable consequences: overtrust in LLMs can lead to security vulnerabilities, data integrity risks, and erroneous decision-making (Thorne, 2024), while undertrust can hinder adoption, reducing the potential benefits of LLMs (Balayn et al., 2024).
@@
The alignment of trust between LLMs and practitioners shapes the range of possible trust relationships in software engineering settings. These relationships involve not only the trustor, the party that decides whether to trust and the trustee, the entity that is trusted (Harjit Sekhon and Devlin, 2014), but also the specific activity in which trust occurs (Balayn et al., 2024). For instance, a practitioner (i.e., a trustor) may choose to accept or reject a code snippet generated by an LLM such as GitHub Copilot (i.e., trustee). In addition, trust perceptions vary according to the trustor, the SE task, and the domain (Balayn et al., 2024). A developer, for example, might confidently rely on an LLM to generate boilerplate code but hesitate to use it for test case generation due to concerns about domain-specific accuracy. Understanding these core trust concepts is essential to unlock the complexities of trust relationships in LLM-assisted software engineering.
@@
However, core concepts such as trust, distrust, and trustworthiness are often used interchangeably, obscuring their precise meanings (Sharma et al., 2024). As a result, fundamental questions about trust in LLM4SE remain unresolved. What factors shape trust and do these factors vary across experience levels, such as between expert and novice developers, or across different SE tasks? Is distrust simply the absence of trust or does it represent a distinct state? Addressing these unresolved questions is crucial for successfully adopting LLMs in SE workflows, as demonstrated by research findings in the field of Human-Computer Interaction (HCI) (Choung et al., 2022; Glikson and Woolley, 2020). These HCI studies have shown that trust is a crucial factor that influences the willingness of stakeholders to adopt LLMs. Unfortunately, existing studies provide a limited foundational understanding of how trust is defined, the factors that influence it, and the metrics used to assess it.
@@
To establish this foundational understanding, we must first assess the current state of research on trust in LLM4SE. Although some existing work has explored aspects of trust in LLMs and even examined factors influencing trust in specific LLM tools within SE, to the best of our knowledge, no comprehensive review has yet been conducted. This absence of a systematic synthesis makes it difficult to determine what aspects of trust have been explored and where significant gaps remain. Previous reviews in broader Artificial Intelligence (AI) contexts provide valuable conceptual insights, but often lack the necessary focus on SE-specific tasks, developer interactions, and the unique challenges posed by LLMs in this domain (Huang et al., 2024; Aljohani et al., 2025; Liu et al., 2024c). Consequently, the absence of a structured overview of trust concepts hinders efforts to define key trust-related concepts, identify influencing factors, and develop reliable evaluation methodologies. Furthermore, the perspective of SE practitioners on the definitions, factors, and metrics of trust remains largely underexplored, resulting in a disconnect between academic research and real-world needs. Without a clear understanding of the existing landscape, advancing trust-aware LLM systems for SE remains a significant challenge.
@@
To address these challenges, we have taken a comprehensive approach that combines a literature review with practitioner insights. First, we review $18$ SE-focused articles to examine how trust-related concepts are defined, what factors influence trust in LLM4SE, and what metrics are used for evaluation. Second, we conducted a survey study with $25$ SE practitioners to capture real-world perspectives on trust in LLM4SE, identify task-specific trust factors, and explore variations in trust perceptions at different levels of expertise. In addition, we carried out a complementary analysis of $70$ articles from disciplines such as Deep Learning (DL), HCI, and Automation to enrich and complement the trust conceptualization of other fields.
@@
By integrating findings from the software engineering literature and practitioner experience, we offer a comprehensive perspective on the current state of trust research in LLM4SE. Beyond identifying gaps in definitions, influencing factors, and evaluation methodologies, our study states the limitations of existing approaches and highlights the urgent need for standardized trust assessment frameworks. Furthermore, we examine the disconnect between academic research and practitioner expectations, revealing key areas where trust calibration – the alignment of user trust level with actual reliability or trustworthiness of a system (Turner et al., 2024) – poses challenges in real-world SE workflows. To improve these aspects, we propose concrete steps to validate trust definitions, improve trust factors, and establish trust metrics. By systematically synthesizing these findings, this work lays the foundations for future advances in trust-based LLM integration, offering concrete directions to improve the reliability, interpretability, and user confidence in these systems.
@@
We structure our investigation around three key Research Questions (RQs) to map current understanding of trust concepts in LLM4SE. Through a systematic review of the literature and a survey study with SE practitioners, we examine how trust, distrust, and trustworthiness are defined in the context of LLM4SE (RQ1), identify the factors that shape trust perceptions (RQ2), and analyze existing approaches to measuring trust in LLM (RQ3). Our findings reveal that while LLMs are increasingly adopted in SE, trust remains a loosely defined and inconsistently measured construct. Existing research neglects domain-specific definitions of trust-related concepts, often borrowing from other disciplines without adaptation to SE. In addition, trust factors, such as precision, interpretability, robustness, and workflow integration, are not consistently considered in different studies, leading to a fragmented understanding of the factors driving trust in LLM4SE. Finally, trust evaluation remains an open challenge, with limited efforts to establish standardized multidimensional metrics that account for the complexity of trust in SE workflows.
@@
    We demonstrate a current understanding of trust definitions, factors that influence trust in LLM4SE, and the metrics used to evaluate trust in LLM4SE, through an extensive review of the literature, including $18$ articles focused on SE, and analysis $70$ articles from the broader trust literature.
@@
    We complement our review with a survey study of $25$ participants to map current research and gather the perspectives of practitioners on trust in LLMs in SE. We highlight the gaps between the literature and practical scenarios and provide a roadmap for future research.
@@
    We propose precise definitions for trust, distrust, and trustworthiness in LLM4SE, integrating insights from both existing literature and practitioners’ perspectives. We also identify key factors influencing trust in LLM4SE, grouping them into model-specific and user-centric factors. Furthermore, we highlight the limitations of single-item metrics and advocate for comprehensive evaluation frameworks that better capture the multifaceted nature of trust in LLM4SE.
@@ 2. Why are Trust, Distrust, and Trustworthiness relevant for researching LLMs in Software Engineering?
Misplaced trust, whether excessive or insufficient, can lead to poor decisions and unintended consequences, such as security vulnerabilities, code smells, and data leaks (Thorne, 2024; Mohsin et al., 2024; Balayn et al., 2024). For example, researchers recently discovered that AI-powered code assistants, including GitHub Copilot and ChatGPT, often generate insecure code snippets that developers often do not recognize as flawed (Sandoval et al., 2023; Pearce et al., 2025). The study also found that when LLM-generated suggestions contained subtle security vulnerabilities, developers were more likely to trust and use them without adequate scrutiny, leading to increased security risks (Pearce et al., 2025). This tendency to over-rely on LLM-generated code without thorough verification goes beyond individual mistakes: it reflects a growing shift in how developers interact with LLMs and make critical coding decisions. As LLMs become integral to software development, they do more than suggest code; they influence workflows, assumptions, and even risk perception. The growing reliance on LLMs in SE requires a structured approach to understanding trust, distrust, and trustworthiness (Huang et al., 2024; Lo, 2023).
@@
Structured understanding of trust in LLM4SE can draw inspiration from research in other fields, such as the ethics of HCI and AI. Studies in these fields highlight key factors that influence trust, including accuracy, interpretability, and transparency (Glikson and Woolley, 2020; Sharma et al., 2024). However, trust in SE is particularly complex due to the technical nature of the domain, the potential for hidden errors in generated code, and the need to align with best practices (Cui et al., 2024; Balayn et al., 2024). Unlike traditional software tools, LLMs produce nondeterministic outputs, which means that the same prompt can produce different responses, making trust assessment even more challenging (Zhu et al., 2024; Ye et al., 2023). Recognizing these challenges, researchers have called for a deeper investigation of trust in SE, e.g., D. Lo (Lo, 2023) emphasizes that as software development moves toward Software 2.0, where LLMs play a central role in coding, ensuring trustworthiness in AI-assisted development is imperative.
@@
Although these calls for research highlight the urgency of understanding trust in LLM4SE, existing studies have not yet fully addressed this issue. Despite the significance of trust in AI-powered SE tools, existing research on LLM4SE has largely overlooked definitions, factors, and metrics of trust. Our review identifies three key challenges related to trust in software engineering research. First challenge ($CH_{1}$) is, only a small fraction of SE studies explicitly define trust, and the distinctions between trust, distrust, and trustworthiness remain underexplored, making it difficult to assess their role in real-world SE workflows. Second challenge ($CH_{2}$), while the broader trust literature provides valuable insight into the factors that influence trust, these perspectives have not been adequately integrated into SE research, leaving a gap in understanding what affects trust in developer interactions with LLMs. Third challenge ($CH_{3}$), there is a noticeable absence of standardized metrics to measure trust, and the perspectives of software practitioners on how trust should be quantified and evaluated are not well represented.
@@
To address these challenges, we conduct a structured literature review, analyze trust research beyond SE, and complement our findings with a practitioner survey. In doing so, our goal is to provide a comprehensive understanding of trust in LLM4SE and lay the groundwork for future research and the development of reliable and trustworthy LLM-powered coding assistants aligned with developers.
@@ 3. Research Questions
This study investigates the trust in LLM4SE through three RQs. Each question is addressed using a systematic review of the trust literature in LLM4SE and a survey with practitioners. To provide additional context, each research question is complemented with an analysis of the underlying trust concept from the broader trust literature.
Figure 1. An Integrated Conceptual Model for Trust in LLMs for Software Engineering
@@ 1. RQ1
    Definition of Trust: How is trust in LLM defined and conceptualized in the context of SE? This question aims to explore how trust, distrust, and trustworthiness are defined by analyzing the existing literature on LLM4SE and the perspectives of practitioners and addressing $CH_{1}$. These findings are contextualized through definitions found in the wider trust literature to highlight key similarities, differences, and gaps. By analyzing all these perspectives, we provide refined, evidence-based definitions that align with academic and practitioner expectations.
@@ 2. RQ2
    Antecedents of Trust: What are the antecedents of trust (i.e., factors that can lead to trust) in LLM in the context of SE? This question investigates the key factors that lead to the formation of trust in LLM4SE, drawing from both the literature of LLM4SE and practitioner experiences and addressing $CH_{2}$. An additional context is provided by analyzing trust factors in established trust models and related domains. The analysis provides actionable recommendations for fostering trust in LLM4SE.
@@ 3. RQ3
    Trust Metrics: How is trust in LLMs measured in the context of SE? This question evaluates the existing metrics used to assess trust in LLM4SE to address $CH_{3}$. We investigate existing trust assessment methods in LLM4SE, examine how professionals evaluate trust in real world settings, and compare these with broader trust evaluation frameworks. By integrating these perspectives, we propose a road map for developing a comprehensive and domain-specific trust metric for LLM4SE.
Fig. 1 presents an overall conceptual framework of our work, illustrating the interrelationships between trust definitions, antecedents, and metrics in LLM-assisted software engineering.